# Smoothed Analysis of Online Metric Matching with a Single Sample: Beyond Metric Distortion

### v1

@@ Abstract
We study this problem in the Euclidean metric $0,1]^{d}$, when servers are adversarial and requests are independently drawn from distinct distributions that satisfy a mild smoothness condition. Our main result is an $O(1)$-competitive algorithm for $d\neq 2$ that requires no distributional knowledge, relying only on a single sample from each request distribution. To our knowledge, this is the first algorithm to achieve an $o(\log n)$ competitive ratio for non-trivial metrics beyond the i.i.d. setting. Our approach bypasses the $\Omega(\log n)$ barrier introduced by probabilistic metric embeddings: instead of analyzing the embedding distortion and the algorithm separately, we directly bound the cost of the algorithm on the target metric of a simple deterministic embedding. We then combine this analysis with lower bounds on the offline optimum for Euclidean metrics, derived via majorization arguments, to obtain our guarantees.
@@ 1 Introduction
When servers and requests are adversarially chosen, [MNP06] give an $O(\log^{3}n)$-competitive algorithm for general metrics, later improved to $O(\log^{2}n)$ [BBGN14]. For Euclidean metrics, an $O(\log n)$-competitive algorithm is given by [GL12]. On the other hand, there exists an $\Omega(\log n)$ competitive-ratio lower bound for uniform metrics [^1] A metric is uniform if every pair of distinct points has distance $1$. [MNP06], which holds even in the random-order arrival model [Rag16]. Moreover, no algorithm can achieve an $o(\sqrt{\log n})$ competitive ratio for line metrics [PS23].
The fully adversarial model is often considered overly pessimistic, since worst-case scenarios rarely arise in practice. A more realistic assumption is that requests are i.i.d. sampled, while servers remain adversarial. In this setting, [GGPW19] give an $O((\log\log\log n)^{2})$-competitive algorithm for general metrics and distributions, and their algorithm is $O(1)$-competitive for tree metrics. Recently, [YY25] show that the above setting further reduces to the case where all servers and requests are i.i.d. sampled from the same distribution. Building on this, they obtain an $O(1)$-competitive algorithm for Euclidean metrics with smooth distributions, where a distribution is smooth if it admits a density with respect to the uniform measure upper bounded by a constant.
The i.i.d. model yields nearly tight guarantees, but it is often unrealistic: in many applications, different requests follow markedly different distributions. A more promising approach is smoothed analysis, which has been widely successful in both theory and practice, and is frequently used to explain the strong empirical performance of heuristics with poor worst-case guarantees [ST09, CGVY24, HRS24, CU25]. In this model, the adversary specifies a smooth distribution over inputs, and the algorithm is required to perform well in expectation. In this paper, we adopt the smoothed analysis framework to study online metric matching and obtain improved competitive ratios under significantly weaker assumptions than those in the i.i.d. setting.
@@ 1.1 Our Results
We focus on the Euclidean metric $0,1]^{d}$, assuming that servers are adversarial and requests are independently drawn from distinct smooth distributions. As our main result, we give an algorithm which requires no knowledge of the distributions and is $O(1)$-competitive for $d\neq 2$, given access to one sample from each of the $n$ request distributions (Theorem 5.1). In particular, the algorithm does not need to know the correspondence between distributions and samples, and its competitive ratio depends polynomially on $d$ and the smoothness parameter, which are usually treated as constants that do not grow with $n$ [GL12, Kan22, YY25]. To the best of our knowledge, this is the first algorithm to achieve an $o(\log n)$ competitive ratio for non-trivial metrics in a setting beyond the i.i.d. assumption, even when the request distributions are fully known to the algorithm (see Table 1 for a thorough comparison with prior work).
@@
Adversarial Adversarial General None $O(\log^{2}n)$ [BBGN14]
Adversarial Adversarial Euclidean None $O(\log n)$ [GL12]
i.i.d. Adversarial General Full knowledge $O((\log\log\log n)^{2})$ [GGPW19]
i.i.d. Adversarial Tree Full knowledge $O(1)$ [GGPW19]
i.i.d. smooth Adversarial Euclidean $O(n^{2})$ samples $O(1)$ for $d\neq 2$ [YY25]
i.i.d. uniform Adversarial Euclidean Full knowledge $O(1)$ [YY25]
@@
Table 1: Comparison with prior work on online metric matching with the state-of-the-art results. Our result is the first to achieve an $o(\log n)$ competitive ratio for nontrivial metrics under distributional assumptions strictly weaker than i.i.d.
@@ Discussion.
Our sample-based assumption—one sample per request distribution—is a natural and practical assumption, widely adopted in recent work on online algorithms [AKW14, RWW20, KNR22, DKL+24, GSW25]. This assumption is also necessary to some extent, as it is unclear how an algorithm could leverage the distributional properties of requests without some information, even in simpler i.i.d. settings. In comparison, our assumptions are weaker than those in prior work for the i.i.d. model: the algorithm of [GGPW19] requires full knowledge of the distribution, and the algorithm of [YY25] needs $O(n^{2})$ samples.
Finally, we turn to the role of the dimension $d$. Two-dimensional space is a critical case for Euclidean matching: prior works highlight that the plane behaves fundamentally differently from both the line and higher dimensions [Tal22, Kan22, YY25]. Consistent with this, our analysis does not yield a constant-competitive bound in $d=2$. Nevertheless, our second algorithm achieves strictly better performance in two dimensions than our main algorithm, offering partial progress. Closing the gap in two dimensions remains an important direction for future research.
@@ 1.2 Technical Overview
Our approach builds on the classical paradigm of metric embeddings but departs in a crucial way. Previous algorithms embed the input metric into a Hierarchically well-Separated Tree (HST) metric and analyze two pieces separately: the distortion of the embedding and the competitive ratio of the algorithm on the HST [MNP06, GL12, BBGN14]. This separation is what forces the $\Omega(\log n)$ barrier. Our key idea is to bypass the distortion step entirely: we analyze the algorithm’s cost directly in the embedding HST metric, using the non-contractivity of the embedding to argue about the original metric. This avoids the logarithmic loss while retaining the algorithmic utility of the HST. With this perspective in place, we now recall the definitions of metric embeddings and HSTs.
@@
A (non-contractive) deterministic embedding of a metric $(\mathcal{X},\delta)$ into another metric $(\mathcal{X}^{\prime},\delta^{\prime})$ is a mapping $f:\mathcal{X}\to\mathcal{X}^{\prime}$ satisfying $\delta(x,y)\leq\delta^{\prime}(f(x),f(y))$ for all $x,y\in\mathcal{X}$, and the distortion of an embedding $f$ is the smallest $\kappa\geq 1$ such that $\delta^{\prime}(f(x),f(y))\leq\kappa\cdot\delta(x,y)$ for all $x,y\in\mathcal{X}$. A probabilistic embedding is a distribution over deterministic embeddings, and the distortion of a probabilistic embedding $f$ is the smallest $\kappa\geq 1$ such that $\mathbb{E}[\delta^{\prime}(f(x),f(y))]\leq\kappa\cdot\delta(x,y)$ for all $x,y\in\mathcal{X}$. It is not hard to see that given an embedding of the input metric to another (usually simpler) metric with distortion $\kappa_{1}$, and an algorithm for the latter metric with competitive ratio $\kappa_{2}$, we get an algorithm for the input metric with competitive ratio $\kappa_{1}\cdot\kappa_{2}$. In other words, a metric embedding with distortion $\kappa$ reduces the problem for a complicated metric to the same problem for a simple metric with the cost of an additional factor $\kappa$ in the competitive ratio.
@@
One of the most popular simple metrics is the HST metric, which is induced by the distance function on an HST. In particular, an $\alpha$-HST for $\alpha\geq 1$ is a rooted tree where all the leaf nodes are at the same depth. Its edge lengths are defined hierarchically: assuming root-adjacent edges have length $1$, and for every internal node $v$, the edge to its parent is $\alpha$ times longer than each edge to its children (see Definition 2.5 for a formal definition and Figure 1 for an illustration). It is known that every $n$-point metric can be randomly embedded into an $\alpha$-HST metric with distortion $O(\alpha\log n)$ [FRT04], which, when used as a reduction for online problems, contributes an $O(\log n)$ factor to the competitive ratio.
@@
To bypass the $O(\log n)$ distortion factor, we avoid the usual two-step analysis that handles the distortion of the metric embedding and the competitive ratio for the HST metric separately. Instead, we bound directly the cost the algorithm incurred on the HST; by the non-contractivity of the embedding, this also upper-bounds the algorithm’s cost in the original metric. We then compare this upper bound to a lower bound for the offline optimum in the original metric, yielding the desired competitive ratio. Notably, a simple deterministic embedding—despite having unbounded worst-case distortion—suffices for our analysis.
@@
In more detail, we use the canonical dyadic partition of $0,1]^{d}$ to define a $2^{d}$-ary $2$-HST metric of height $h$ (a tunable parameter we optimize later), where an HST is $\Delta$-ary if every internal node has exactly $\Delta$ children. To describe the embedding, the root of the HST corresponds to $0,1]^{d}$, each node is partitioned into $2^{d}$ subcubes with half the side length, and leaves at depth $h$ correspond to disjoint subcubes of side-length $2^{-h}$ (see Figure 2 for an illustration). Then, we map each point $x\in[0,1]^{d}$ to the unique leaf node whose corresponding cube contains $x$.
@@
Equipped with the metric embedding, the problem reduces to analyzing HST metrics, where we build on two classic algorithms: the Random-Subtree algorithm of [GL12] and the algorithm of [BBGN14]. Our key observation is that the expected cost analysis for both algorithms reduces to bounding the fluctuations in the number of requests in each subtree. By controlling these fluctuations (via standard deviations of Poisson–binomial counts) and applying concavity, we show that the worst case occurs when requests are uniformly distributed across children. This yields distribution-free upper bounds for the cost (Theorems 3.1 and 4.1), so no smoothness assumption is needed for the upper bounds.
@@
Smoothness enters only in the lower bound on the offline optimum (Lemma 5.3). For $d\geq 2$, the bound follows directly from standard nearest-neighbor distance estimates. The proof for $d=1$ is subtler. Any imbalance between the numbers of servers and requests in a subinterval of $[0,1]$ forces a proportional number of matches to cross the subinterval’s boundary, contributing to the total cost. Inspecting all subintervals of length $L\in(0,1)$ yields a lower bound for the offline optimum, which we refer to as obstacle to matching at length scale $L$. It is known that, when all servers and requests are uniformly distributed, the largest obstacle to matching for $d=1$ occurs at the scale of a constant length [Kan22]. To generalize this reasoning to the case where requests are drawn from distinct smooth distributions, we derive a lower bound for the obstacle to matching at a length scale proportional to the smoothness parameter. Using majorization and concavity arguments, we show that this lower bound is minimized when requests are as concentrated as possible. The desired lower bound then follows from the anti-concentration properties of smooth distributions.
@@
Finally, we apply the result in [YY25] to incorporate the provided samples from the request distributions in a black-box manner (Lemma 2.4). Combining the above upper bounds for the algorithms with the lower bounds for the offline optimum yields our main result.
@@ 1.3 Related Work
For adversarial servers and requests, [KP93, KMV94, Rag16] give (2n1)(2n-1)-competitive deterministic algorithms for general metrics, which is optimal. [Rag16] provides a primal-dual deterministic algorithm that is $O(\log n)$-competitive for general metrics in the random-order arrival model, which is later shown to admit near instance-optimality [NR17] and an $O(\log n)$ competitive ratio for line metrics [Rag18].
When all servers and requests are uniformly distributed on the Euclidean metric, [Kan22] gives an $O(1)$-competitive algorithm, which applies the same deterministic metric embedding as ours. Nevertheless, they employ a different algorithm for the HST metric, and it is unclear how to generalize their analysis to non-identical request distributions. When all servers and requests are i.i.d. drawn from a general distribution, [CKKZ23] present an algorithm for Euclidean metrics with nearly optimal regret, which is defined as the difference between the cost of the algorithm and the offline optimum. [BFP23] show that the Greedy algorithm is $O(1)$-competitive for line metrics when all servers and requests are uniformly distributed.
[AALS22] initiate the study of unbalanced markets, where servers outnumber requests by a constant factor, and they show that Greedy is $O(\log^{3}n)$-competitive when all servers and requests are uniformly distributed on a line, which is subsequently improved to $O(1)$ [BFP23]. [Kan22] gives an $O(1)$-competitive algorithm for unbalanced markets when all servers and requests are uniformly distributed on the Euclidean metric. When servers are adversarial and requests are i.i.d. drawn from a smooth distribution, [YY25] achieve competitive-ratio and regret guarantees for unbalanced markets and Euclidean metrics.
Several variants of online metric matching have also been studied, which include online transportation [HI25, AMPU25], online metric matching with recourse [GKS20, MN25], online min-cost perfect matching with delays [EKW16, AAC+17], online metric matching with stochastic arrivals and departures [AS22, AAS25], online min-weight perfect matching [BFT24], and online matching in geometric random graphs [SNL+23].
@@
For smoothed analysis of other online problems, prior studies primarily focus on online learning [KMR+18, HHSY22, BDGR22, DHZ23, HRS24] and online discrepancy minimization [BJM+22a, BJM+22b, HRS24], whose goal is to minimize regret. Regarding smoothed competitive analysis, [BLM+06] consider the multi-level feedback algorithm for non-clairvoyant scheduling, and [SS05] analyze the work function algorithm for metrical task systems. More recently, [CU25] conduct smoothed analysis of classic online metric problems including $k$-server, $k$-taxi, and chasing small sets, where requests are drawn from smooth distributions, and they achieve significantly improved competitive ratios compared to the adversarial setting. In particular, they allow the distribution followed by each request to depend on the realizations of the past requests and the decisions made by the algorithm thus far, and their algorithms require no distributional knowledge. However, their technique does not results in improved algorithms for online metric matching.
@@
A growing line of work studies online algorithms that receive samples from the underlying distributions to go beyond the strong assumption of knowing the distributions in full. Pioneered by [AKW14], competitive guarantees are achieved for sample-based prophet inequalities under various combinations of the arrival model, the number of samples, and combinatorial constraints [CDFS19, RWW20, CDF+22, CCES24, CZ24, FLT+24]. This paradigm has also been adopted by literature on online resource allocation [DKL+24, GSW25] and online weighted matching [KNR22].

### v2

@@ Abstract
We study this problem in $0,1]^{d}$ with the Euclidean metric, when servers are adversarial and requests are independently drawn from distinct distributions that satisfy a mild smoothness condition. Our main result is an $O(1)$-competitive algorithm for $d\neq 2$ that requires no distributional knowledge, relying only on a single sample from each request distribution. To our knowledge, this is the first algorithm to achieve an $o(\log n)$ competitive ratio for non-trivial metrics beyond the i.i.d. setting. Our approach bypasses the $\Omega(\log n)$ barrier introduced by probabilistic metric embeddings: instead of analyzing the embedding distortion and the algorithm separately, we directly bound the cost of the algorithm on the target metric space of a simple deterministic embedding. We then combine this analysis with lower bounds on the offline optimum for Euclidean metrics, derived via majorization arguments, to obtain our guarantees.
@@ 1 Introduction
When servers and requests are adversarially chosen, [DBLP:conf/soda/MeyersonNP06] give an $O(\log^{3}n)$-competitive algorithm for general metrics, later improved to $O(\log^{2}n)$ [DBLP:journals/algorithmica/BansalBGN14]. For Euclidean metrics, an $O(\log n)$-competitive algorithm is given by [DBLP:conf/icalp/GuptaL12]. On the other hand, there exists an $\Omega(\log n)$ competitive-ratio lower bound for uniform metrics [^1] A metric is uniform if every pair of distinct points has distance $1$. [DBLP:conf/soda/MeyersonNP06], which holds even in the random-order arrival model [DBLP:conf/approx/Raghvendra16]. Moreover, no algorithm can achieve an $o(\sqrt{\log n})$ competitive ratio for line metrics [DBLP:journals/talg/PesericoS23].
The fully adversarial model is often considered overly pessimistic, since worst-case scenarios rarely arise in practice. A more realistic assumption is that requests are i.i.d. sampled, while servers remain adversarially chosen. In this setting, [DBLP:conf/icalp/GuptaGPW19] give an $O((\log\log\log n)^{2})$-competitive algorithm for general metrics and distributions, and their algorithm is $O(1)$-competitive for tree metrics. Recently, [yang2025online] show that the above setting further reduces to the case where all servers and requests are i.i.d. sampled from the same distribution. Building on this, they obtain an $O(1)$-competitive algorithm for Euclidean metrics with smooth distributions, where a distribution is smooth if it admits a density with respect to the uniform measure upper bounded by a constant.
The i.i.d. model yields nearly tight guarantees, but it is often unrealistic: in many applications, different requests follow markedly different distributions. A more promising approach is smoothed analysis, which has been widely successful in both theory and practice, and is frequently used to explain the strong empirical performance of heuristics with poor worst-case guarantees [DBLP:journals/cacm/SpielmanT09, DBLP:conf/soda/ChenGVY24, DBLP:conf/stoc/Anastos0M25]. In this model, the adversary first selects an adversarial input, which is then randomly perturbed by nature via adding a small noise, and the algorithm is required to perform well in expectation. Since the perturbed input always follows a smooth distribution, in the more modern and general formulation of smoothed analysis, the adversary directly specifies a smooth distribution over inputs [DBLP:journals/jacm/HaghtalabRS24, DBLP:conf/esa/CoesterU25]. In this paper, we adopt the smoothed analysis framework to study online metric matching and obtain improved competitive ratios under significantly weaker assumptions than those in the i.i.d. setting.
@@ 1.1 Our Results
We focus on $0,1]^{d}$ with the Euclidean metric, assuming that servers are adversarial and requests are independently drawn from distinct smooth distributions. As our main result, we give an algorithm which requires no knowledge of the distributions and is $O(1)$-competitive for $d\neq 2$, given access to one sample from each of the $n$ request distributions (Theorem 5.1). In particular, the algorithm does not need to know the correspondence between distributions and samples, and its competitive ratio depends polynomially on $d$ and the smoothness parameter, which are usually treated as constants that do not grow with $n$ [DBLP:conf/icalp/GuptaL12, DBLP:conf/sigecom/Kanoria22, yang2025online]. To the best of our knowledge, this is the first algorithm to achieve an $o(\log n)$ competitive ratio for non-trivial metrics in a setting beyond the i.i.d. assumption, even when the request distributions are fully known to the algorithm (see Table 1 for a thorough comparison with prior work).
@@
Adversarial Adversarial General None $O(\log^{2}n)$ [DBLP:journals/algorithmica/BansalBGN14]
Adversarial Adversarial Euclidean None $O(\log n)$ [DBLP:conf/icalp/GuptaL12]
i.i.d. Adversarial General Full knowledge $O((\log\log\log n)^{2})$ [DBLP:conf/icalp/GuptaGPW19]
i.i.d. Adversarial Tree Full knowledge $O(1)$ [DBLP:conf/icalp/GuptaGPW19]
i.i.d. smooth Adversarial Euclidean $O(n^{2})$ samples $O(1)$ for $d\neq 2$ [yang2025online]
i.i.d. uniform Adversarial Euclidean Full knowledge $O(1)$ [yang2025online]
@@
Table 1: Comparison with prior work on online metric matching with the state-of-the-art results. Our result is the first to achieve an $o(\log n)$ competitive ratio for non-trivial metrics under distributional assumptions strictly weaker than i.i.d.
@@ Discussion.
Our sample-based assumption—one sample per request distribution—is a natural and practical assumption, widely adopted in recent work on online algorithms [DBLP:conf/soda/AzarKW14, DBLP:conf/innovations/RubinsteinWW20, DBLP:conf/soda/KaplanNR22, DBLP:conf/focs/DuttingKLR024, DBLP:conf/stoc/Ghuge0W25]. This assumption is also necessary to some extent, as it is unclear how an algorithm could leverage the distributional properties of requests without some information, even in simpler i.i.d. settings. In comparison, our assumptions are weaker than those in prior work for the i.i.d. model: the algorithm of [DBLP:conf/icalp/GuptaGPW19] requires full knowledge of the distribution, and the algorithm of [yang2025online] needs $O(n^{2})$ samples.
Finally, we turn to the role of the dimension $d$. Two-dimensional space is a critical case for Euclidean matching: prior works highlight that the plane behaves fundamentally differently from both the line and higher dimensions [talagrand2022upper, DBLP:conf/sigecom/Kanoria22, yang2025online]. Consistent with this, our analysis does not yield a constant-competitive bound in $d=2$. Nevertheless, our second algorithm achieves strictly better performance in two dimensions than our main algorithm, offering partial progress. Closing the gap in two dimensions remains an important direction for future research.
@@ 1.2 Technical Overview
Our approach builds on the classical paradigm of metric embeddings but departs in a crucial way. Previous algorithms embed the input metric space into Hierarchically well-Separated Trees (HST) and analyze two pieces separately: the distortion of the embedding and the competitive ratio of the algorithm on HSTs [DBLP:conf/soda/MeyersonNP06, DBLP:conf/icalp/GuptaL12, DBLP:journals/algorithmica/BansalBGN14]. This separation is what forces the $\Omega(\log n)$ barrier. Our key idea is to bypass the distortion step entirely: we analyze the algorithm’s cost directly in the resulting HST, using the non-contractivity of the embedding to argue about the original metric space. This avoids the logarithmic loss while retaining the algorithmic utility of the HST. With this perspective in place, we now recall the definitions of metric embeddings and HSTs.
@@
A (non-contractive) deterministic embedding of a metric space $(\mathcal{X},\delta)$ into another metric space $(\mathcal{X}^{\prime},\delta^{\prime})$ is a mapping $f:\mathcal{X}\to\mathcal{X}^{\prime}$ satisfying $\delta(x,y)\leq\delta^{\prime}(f(x),f(y))$ for all $x,y\in\mathcal{X}$, and the distortion of an embedding $f$ is the smallest $\kappa\geq 1$ such that $\delta^{\prime}(f(x),f(y))\leq\kappa\cdot\delta(x,y)$ for all $x,y\in\mathcal{X}$. A probabilistic embedding is a distribution over deterministic embeddings, and the distortion of a probabilistic embedding $f$ is the smallest $\kappa\geq 1$ such that $\mathbb{E}[\delta^{\prime}(f(x),f(y))]\leq\kappa\cdot\delta(x,y)$ for all $x,y\in\mathcal{X}$. It is not hard to see that given an embedding of the input metric space to another (usually simpler) metric space with distortion $\kappa_{1}$, and an algorithm for the latter metric space with competitive ratio $\kappa_{2}$, we get an algorithm for the input metric space with competitive ratio $\kappa_{1}\cdot\kappa_{2}$. In other words, a metric embedding with distortion $\kappa$ reduces the problem for a complicated metric to the same problem for a simple metric with the cost of an additional factor $\kappa$ in the competitive ratio.
@@
One of the most popular simple metrics is the HST metric, which is induced by the distance function on an HST. In particular, an $\alpha$-HST for $\alpha\geq 1$ is a rooted tree where all the leaf nodes are at the same depth. Its edge lengths are defined hierarchically: assuming root-adjacent edges have length $1$, and for every internal node $v$, the edge to its parent is $\alpha$ times longer than each edge to its children (see Definition 2.6 for a formal definition and Figure 1 for an illustration). It is known that every $n$-point metric space can be randomly embedded into an $\alpha$-HST with distortion $O(\alpha\log n)$ [DBLP:journals/jcss/FakcharoenpholRT04], which, when used as a reduction for online problems, contributes an $O(\log n)$ factor to the competitive ratio.
@@
To bypass the $O(\log n)$ distortion factor, we avoid the usual two-step analysis that handles the distortion of the metric embedding and the competitive ratio for HSTs separately. Instead, we bound directly the cost the algorithm incurred on the resulting HST; by the non-contractivity of the embedding, this also upper-bounds the algorithm’s cost in the original metric space. We then compare this upper bound to a lower bound for the offline optimum in the original metric space, yielding the desired competitive ratio. Notably, a simple deterministic embedding—despite having unbounded worst-case distortion—suffices for our analysis.
@@
In more detail, we use the canonical dyadic partition of $0,1]^{d}$ to define a $2^{d}$-ary $2$-HST of height $h$ (a tunable parameter we optimize later), where an HST is $\Delta$-ary if every internal node has exactly $\Delta$ children. To describe the embedding, the root of the HST corresponds to $0,1]^{d}$, each node is partitioned into $2^{d}$ subcubes with half the side-length, and the leaves at depth $h$ correspond to disjoint subcubes of side-length $2^{-h}$ (see Figure 2 for an illustration). Then, we map each point $x\in[0,1]^{d}$ to the unique leaf node whose corresponding cube contains $x$.
@@
Equipped with the metric embedding, the problem reduces to designing algorithms for HST metrics, where we adopt the Random-Subtree algorithm of [DBLP:conf/icalp/GuptaL12] and the algorithm of [DBLP:journals/algorithmica/BansalBGN14]. Our key observation is that the expected cost analysis for both algorithms reduces to bounding the fluctuations in the number of requests in each subtree. By controlling these fluctuations (via standard deviations of Poisson–binomial counts) and applying concavity, we show that the worst case occurs when requests are uniformly distributed across children. This yields distribution-free upper bounds for the cost (Theorems 3.1 and 4.1), so no smoothness assumption is needed for the upper bounds.
@@
Smoothness enters only in the lower bound on the offline optimum (Lemma 5.3). For $d\geq 2$, the bound follows directly from standard nearest-neighbor distance estimates. The proof for $d=1$ is subtler. Any imbalance between the numbers of servers and requests in a subinterval of $[0,1]$ forces a proportional number of matches to cross the subinterval’s boundary, contributing to the total cost. Inspecting all subintervals of length $L\in(0,1)$ yields a lower bound for the offline optimum, which we refer to as obstacle to matching at length scale $L$. It is known that, when all servers and requests are uniformly distributed, the largest obstacle to matching for $d=1$ occurs at the scale of a constant length [DBLP:conf/sigecom/Kanoria22]. To generalize this reasoning to the case where requests are drawn from distinct smooth distributions, we derive a lower bound for the obstacle to matching at a length scale proportional to the smoothness parameter. Using majorization and concavity arguments, we show that this lower bound is minimized when requests are as concentrated as possible. The desired lower bound then follows from the anti-concentration properties of smooth distributions.
@@
Finally, we apply the result in [yang2025online] to incorporate the provided samples from the request distributions in a black-box manner (Lemma 2.5). Combining the above upper bounds for the algorithms with the lower bounds for the offline optimum yields our main result.
@@ 1.3 Related Work
For adversarial servers and requests, [DBLP:journals/jal/KalyanasundaramP93, DBLP:journals/tcs/KhullerMV94, DBLP:conf/approx/Raghvendra16] give (2n1)(2n-1)-competitive deterministic algorithms for general metrics, which is optimal. [DBLP:conf/approx/Raghvendra16] provides a primal-dual deterministic algorithm that is $O(\log n)$-competitive for general metrics in the random-order arrival model, which is later shown to exhibit near instance-optimality [DBLP:conf/focs/NayyarR17] and an $O(\log n)$ competitive ratio for line metrics [DBLP:conf/compgeom/Raghvendra18].
When all servers and requests are uniformly distributed on the Euclidean space, [DBLP:conf/sigecom/Kanoria22] gives an $O(1)$-competitive algorithm, which applies the same deterministic metric embedding as ours. Nevertheless, they employ a different algorithm for the HST metric, and it is unclear how to generalize their analysis to non-identical request distributions. When all servers and requests are i.i.d. drawn from a general distribution, [DBLP:conf/sigecom/ChenKKZ23] present an algorithm for Euclidean metrics with nearly optimal regret, which is defined as the difference between the cost of the algorithm and the offline optimum. [DBLP:conf/sigecom/BalkanskiFP23] show that the Greedy algorithm is $O(1)$-competitive for line metrics when all servers and requests are uniformly distributed.
[DBLP:conf/sigecom/AkbarpourALS22] initiate the study of unbalanced markets, where servers outnumber requests by a constant factor, and they show that Greedy is $O(\log^{3}n)$-competitive when all servers and requests are uniformly distributed on a line, which is subsequently improved to $O(1)$ [DBLP:conf/sigecom/BalkanskiFP23]. [DBLP:conf/sigecom/Kanoria22] gives an $O(1)$-competitive algorithm for unbalanced markets when all servers and requests are uniformly distributed on the Euclidean space. When servers are adversarial and requests are i.i.d. drawn from a smooth distribution, [yang2025online] achieve competitive-ratio and regret guarantees for unbalanced markets and Euclidean metrics.
Several variants of online metric matching have also been studied, which include online transportation [DBLP:conf/icalp/HaradaI25, arndt2025competitive], online metric matching with recourse [DBLP:conf/approx/0004KS20, DBLP:journals/algorithmica/MegowN25], online min-cost perfect matching with delays [DBLP:conf/stoc/EmekKW16, DBLP:conf/approx/AshlagiACCGKMWW17], online metric matching with stochastic arrivals and departures [DBLP:journals/ior/AouadS22, DBLP:conf/stoc/AmaniHamedaniAS25], online min-weight perfect matching [DBLP:conf/soda/BhoreFT24], and online matching in geometric random graphs [DBLP:journals/corr/abs-2306-07891].
@@
For smoothed analysis of other online problems, prior studies primarily focus on online learning [DBLP:conf/nips/KannanMRWW18, DBLP:conf/nips/HaghtalabHSY22, DBLP:conf/colt/BlockDGR22, DBLP:conf/sigecom/DurvasulaHZ23, DBLP:journals/jacm/HaghtalabRS24] and online discrepancy minimization [DBLP:conf/innovations/BansalJM0S22, DBLP:conf/icalp/0001JM0S22, DBLP:journals/jacm/HaghtalabRS24], whose goal is to minimize regret. Regarding smoothed competitive analysis, [DBLP:journals/mor/BecchettiLMSV06] consider the multi-level feedback algorithm for non-clairvoyant scheduling, and [DBLP:journals/tcs/SchaferS05] analyze the work function algorithm for metrical task systems. More recently, [DBLP:conf/esa/CoesterU25] conduct smoothed analysis of classic online metric problems including $k$-server, $k$-taxi, and chasing small sets, where requests are drawn from smooth distributions, and they achieve significantly improved competitive ratios compared to the adversarial setting. In particular, they allow the distribution followed by each request to depend on the realizations of the past requests and the decisions made by the algorithm thus far, and their algorithms require no distributional knowledge. However, their technique does not results in improved algorithms for online metric matching.
@@
A growing line of work studies online algorithms that receive samples from the underlying distributions to go beyond the strong assumption of knowing the distributions in full. Pioneered by [DBLP:conf/soda/AzarKW14], competitive guarantees are achieved for sample-based prophet inequalities under various combinations of the arrival model, the number of samples, and combinatorial constraints [DBLP:conf/ec/CorreaDFS19, DBLP:conf/innovations/RubinsteinWW20, DBLP:conf/soda/CaramanisDFFLLP22, DBLP:journals/mor/0001C0S24, DBLP:conf/stoc/CristiZ24, DBLP:conf/sigecom/0001LTWW024]. This paradigm has also been adopted by literature on online resource allocation [DBLP:conf/focs/DuttingKLR024, DBLP:conf/stoc/Ghuge0W25] and online weighted matching [DBLP:conf/soda/KaplanNR22].